Tech Staff Summary 5/10/00

Current Priorities in addition to routine tasks

8 of 12 stations are operational.
Prison at Tiptonville is likely. Smalley is handling the paperwork.
Couple possibilities for Covington. Ellis and Mac are looking for a site.
Couple possibilities for Troy/Obian. Ellis and Mac are looking for a site.
TN river site is a big problem. Just aren't any known good areas that meet specs. Mac is continuing scouting as time permits.

Greg and Steve are organizing a robust test of the two aironets.
Greg is delevoping a scheme for getting the data from the unnamed tower into the computer room without going through the CERI router in the long building.
Mitch will continue to hound the lawyers to get the PPWT agreement and then send to others.
Greg will purchase necessary equipment using constructed equipment code Sheila set up.
Mitch is still waiting on coordinates from Greg of tower sites.

John is plugging away at permissions but the redtape is going slow.
John has hired an assistant.
MTTN node may be delayed until we get some stations up and running.
SWET was a great success. Data are being recorded at CERI. Mitch is looking into the helicorder bug that causes improper sample rates in the gif images.
Possible areas for the other CMG3 include north central AL or central TN/KY area. Mitch is figuring this out and will interface with Herrmann and Buland.
Knoxville/TVA node is still slated for September. Mitch is working with Williams and Benz to get this organized.
Steve is looking into possibilities for getting ATHN continuously.
LRNC has extremely poor telemetry and marginal internet bandwidth. Possible fix for the telemetry is reactivation of High Peak as a repeater. Greg will get with John about this.

Board Shop
The high speed router demo was impressive and we will likely use this in the future. Greg will keep contact with the Salesman in the event a cost effective used system becomes available. Purchase of a new system will not take place at this time.
Current OSHA problems will need to be adressed asap in order to avoid delays on other projects. Greg will look into and arrange to have an operational eyewash station installed, remove the inoperable shower, repair ceiling, and install ventilation.
Greg attended the Hazmat course and arranged for a CERI inspection. Environmental concerns with the boardshop were unfounded. This was confirmed by the UofM ES&H.

Chris took the AC course so we should now be able to maintain our own systems at the nodes.

Ready to move. Due to delays in the DBB and MTTN projects this will get bumped up.
Mitch still needs to complete the code to handle the gain channel. Subroutines are done and offline PII style is available. Need some real ISIS II to experiment with. On the fly will need less robust than offline to avoid potentially large and time intensive buffering.

Steve is still working on the rapid map generation. When complete, this can be attached to the auto email messages and to the newly installed recenteqs system.

Greg will fix the old clock to supply IRIGH.
Mitch has obtained and tested National Instruments 16-bit a2d boards.
Steve has obtained a PC and is getting the mux.
On schedule for 7/31 installation. We will digitize the IRIGH and stamp with pc system time. Running xntp will help keep this accurate. Consider replacing ISDN with DSL.

Paul is set for the test deployment the week of 5/31.

Second tier priorities

Cannabalize PI to make as much as PII operational.
Check with Arch or Jer-Ming for account numbers to buy supplies.
Not sure where the space exists to work on this.

Analog backbone reconfigure is waiting on the lawyers.

EBZ/PWLA telemetry. Another repeater?

LNXT node Secure propane tank.
Secure racks.
replace coax jumpers
replace rg-58 unshielded jumpers

MKTA node

NMAD node

Need to record the SOH channel from the gps clocks to know when sync is lost (they will continue to produce IRIGE based on internal clock).

Backburner priorities.

Station book. Mitch and Steve will install earthworm v5 in June. This will include the full blown schema capable of creating seed.

Status channels. Currently supplied on last 8 channels.

LNXT node permit/install L7 and L8

MKTA node permit/install remaining 3 ISIS

NMAD node